When the first equation is solved we obtain the value of x as 3, when 3 is substituted in the second equation, its value is 10, hence only option A is correct.
4(3x - 3) = 24 or 12x - 12 = 24 or 12x = 36 or x = 3
5x - 5 = 5 * 3 - 5 = 15 - 5 = 10

Please show me how to solve the question
step by step
Thank you

Answer by richwmiller(17219)

Please submit the whole problem with all options.
Just the problem first and then your attempts at solutions.
Is the problem
4*(3x-3)=y or 4*(3x-3)=24
5x-5=y or 5x-5=5